### v2

# Explanation User Interfaces: A Systematic Literature Review

Abstract.
Artificial Intelligence (AI) is one of the major technological advancements of this century, bearing incredible potential for users through AI-powered applications and tools in numerous domains. Being often black-box (i.e., its decision-making process is unintelligible), developers typically resort to eXplainable Artificial Intelligence (XAI) techniques to interpret the behaviour of AI models to produce systems that are transparent, fair, reliable, and trustworthy. However, presenting explanations to the user is not trivial and is often left as a secondary aspect of the system’s design process, leading to AI systems that are not useful to end-users. This paper presents a Systematic Literature Review on Explanation User Interfaces (XUIs) to gain a deeper understanding of the solutions and design guidelines employed in the academic literature to effectively present explanations to users. To improve the contribution and real-world impact of this survey, we also present a platform to support Human-cEnteRed developMent of Explainable user interfaceS (HERMES) and guide practitioners and scholars in the design and evaluation of XUIs.
Explainable Artificial Intelligence (AI), XUI, design, evaluation, visualization, explanation
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: CSURjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: General and reference Surveys and overviewsccs: Human-centered computing User interface designccs: Computing methodologies Artificial intelligence
HCXAI
Human-Centered eXplainable Artificial Intelligence
GDPR
General Data Protection Regulations
LLM
Large Language Model
UCD
User-Centered Design
ML
Machine Learning
DL
Deep Learning
XUI
Explanation User Interface
HCAI
Human-Centered Artificial Intelligence
XAI
eXplainable Artificial Intelligence
AI
Artificial Intelligence
HCD
Human-Centered Design
HCI
Human-Computer Interaction
UI
User Interface
SLR
Systematic Literature Review
EU
European Union
LIME
Locally Interpretable Model-agnostic Explanations
SHAP
SHapley Additive exPlanation
CAM
Class Activation Mapping
Grad-CAM
Gradient-weighted Class Activation Mapping
PDP
Partial Dependence Plot
DiCE
Diverse Counterfactual Explanations
DARPA
Defense Advanced Research Projects Agency
HERMES
a framework for Human-cEnteRed developMent of Explainable user interfaceS
1. Introduction
Recent years have witnessed a remarkable interest in the field of Artificial Intelligence (AI), which has become a prominent topic in both academia and industry. The rapid advancements in AI technologies have led to their widespread adoption across various sectors, including healthcare, finance, and transportation. The Machine Learning subfield of AI has gained particular attention due to its ability to provide discriminative and predictive capabilities, enabling the development of sophisticated applications such as image recognition, natural language processing, and autonomous systems. The increasing complexity of Machine Learning models has raised concerns about their interpretability and transparency, particularly for those models based on deep learning architectures. These models often operate as “black boxes,” making it challenging for users to understand their decision-making processes (Guidotti et al., 2019). This has led to a growing demand for eXplainable Artificial Intelligence (XAI) techniques that can elucidate the decision-making processes of these models. Black-box Machine Learning models pose various threats to their users. For example, the risk of unfair and unjust bias increases when the model’s inner workings are not clear (Mehrabi et al., 2022). Therefore, worldwide legislative bodies are attempting to regulate the use of such systems. For example, the European Union (EU) recently released the AI act with this goal (European Parliament and Council of the European Union, 2024). Most of these regulations advocate the need for AI systems to be accountable, fair, and explainable (European Parliament and Council of the European Union, 2024). In this context, XAI may be beneficial. XAI is a sub-field of AI, initially formalised by the Defense Advanced Research Projects Agency (DARPA) in 2017 (Gunning and Aha, 2019), which focuses on “opening” black-box models, making their inner workings clear, explainable, and interpretable (Guidotti et al., 2019; Gunning and Aha, 2019). Although different techniques for explaining black-box models exist (as will be detailed in Section 2.1), a common and precise definition of “good” explanations has yet to be established. Several studies have pointed out that the majority of the work on XAI is based on researchers’ intuition of what qualifies as a “good” explanation (Miller, 2019; Abdul et al., 2018; Liao and Varshney, 2021; Lipton, 2018; Cheng et al., 2019), effectively framing XAI primarily as an algorithmic problem. As a result, there is a gap between XAI algorithms designed by researchers and their deployment in real-word scenarios  (Liao and Varshney, 2021). To fill this gap, Human-Computer Interaction (HCI) research has recently begun to focus on AI-enabled applications and XAI, as testified by the birth of the research field of Human-Centered Artificial Intelligence (HCAI), which reframes AI-development in a human-centred perspective by focusing on users’ needs (Shneiderman, 2022). Since the first conceptualisation of XAI by DARPA, the design of XAI systems followed two different stages: the design of the algorithmic model itself, and the design of the User Interface (UI) (Chromik and Butz, 2021). Such UIs are usually referred to as Explanation User Interfaces (Chromik and Butz, 2021). As with other types of UIs, XUIs must be designed following a human-centred approach to be successful and usable (ISO, 2018a).
In this study, we present a Systematic Literature Review (SLR), following Kitchenham’s procedure (Kitchenham, 2004), that aims at providing a comprehensive overview of the design of XUIs. Our overall goal is to identify design trends among human-centred XUIs and XAI algorithms to provide a cohesive understanding of the Human-Centered eXplainable Artificial Intelligence (HCXAI) landscape. More precisely, we contribute to the state-of-the-art systematising the knowledge on four main interrelated aspects of XUIs design; (i) The XAI models used to generate the explanations, (ii) The solutions adopted to provide the explanations, (iii) The user studies and human-centred techniques used to evaluate the designs, (iv) The design guidelines and frameworks that can help designers and practitioners. This article is structured as follows: Section 2 presents key background concepts; Section 3 presents the methodology used to conduct the SLR and the four research questions that guided it; Sections 4.1, 4.2, 4.3 and 4.4report the results to answer each of the four research questions; Section 5 presents a set of conclusions drawn by analyzing the dimensions of each RQ; moreover, it presents HERMES, a platform to guide practitioners in the design and evaluation of XUIs; Section 6 reports open research challenges; Section 7 presents the limitations of our study, highlighting any potential threat to its validity; finally, Section 8 concludes the article.
2. Background
This section provides an overview of the background of our work. It provides an overview of the field of XAI, reprising a classification of methods already available in the literature that can guide in understanding the results of this SLR. Furthermore, it explores the need for a human-centred process and its implications for XUIs.
2.1. Explainable Artificial Intelligence
[…]
Interpretable models:
A small set of interpretable models, such as decision trees and decision rules, can be used to explain more complex models by approximating their global or local behaviour. For example, a decision tree can approximate the behaviour of a complex model, allowing its interpretation as a chain of if-then-else rules. This line of research builds upon foundational work in extracting symbolic rules from trained artificial neural networks (Andrews et al., 1995) and neuro-fuzzy systems (Mitra and Hayashi, 2000), recently systematized in reviews on symbolic knowledge extraction and injection (Ciatto et al., 2024).
Features Importance:
Explanations may be provided by describing each feature’s weight (i.e., importance) in the decision-making process. Such techniques are referred to as “posthoc” since they can be employed after the complex model has been built. An example of such techniques is Locally Interpretable Model-agnostic Explanations (LIME) (Ribeiro et al., 2016), a technique to explain individual predictions by approximating the local behavior of a model with a simpler, interpretable model around a specific instance.
Shapley Values:
A particular instance of explanations based on feature importance is SHapley Additive exPlanation (SHAP), a technique based on cooperative game theory which uses Shapley values to attribute a model’s output to each input feature, ensuring consistency and local accuracy in explanations (Lundberg and Lee, 2017). Compared to an approach like LIME, which can only provide local explanations, SHAP can also produce global explanations, at the cost of a higher computational expense.
Saliency Masks:
Saliency masks highlight subsets of the original instance (e.g., set of pixels of an image or words in a sentence) mainly responsible for a certain outcome. Thus, a “salient mask” visually highlights the aspects of the input instance determining the outcome. Saliency masks are generally used to explain deep neural networks. E.g., Class Activation Mapping (CAM) (Zhou et al., 2016), and its generalisation Gradient-weighted Class Activation Mapping (Grad-CAM) (Selvaraju et al., 2020), highlight the contents of an image based on an attention mechanism, which is visualised as an explanation. Saliency masks are also helpful for text: rationales (Lei et al., 2016) are a short and coherent piece of text (e.g., a sentence) that represents a sufficient subset of words that can be used to predict the original text on their own.
Sensitivity Analysis:
This technique can be used to inspect black-box models by evaluating the correlation between the uncertainties of their inputs and outputs (Saltelli, 2002). Sensitivity analysis is generally used to develop visual tools for inspecting black boxes.
Partial Dependence Plot:
Partial Dependence Plots are graphical representations showing the relationship between one or more input features and the model prediction while averaging the remaining features’ effects. PDPs help understand how changes in certain input features affect the output of a black-box model. Individual Conditional Expectation (Goldstein et al., 2015) is an extension of PDPs that allows for visualising the relationship between a feature and the prediction outcome, considering every individual instance in the dataset, instead of averaging over all instances.
Exemplar-based/Prototype-based Explanations:
A prototype (or archetype) is an object that is representative of similar instances (such as instances belonging to a certain class). Together with the outcome, these explanations return a prototype, which is very similar to the classified record, to highlight the rationale behind the prediction. Prototypes can be obtained, e.g., as a result of the averages of the features of a set of points (Fouché and Langit, 2011).
Counterfactual Explanations:
Counterfactual explanations illustrate how minimal changes to input features can alter the model’s prediction. They answer “what-if” questions by providing examples of how the input instance needs to be changed into a similar one to arrive at a different class in the output. Diverse Counterfactual Explanations (DiCE) is a method that generates multiple diverse counterfactual instances for a given instance to help users understand different ways to achieve a desired outcome (Mothilal et al., 2020).
Neurons Activation:
These interpretability techniques investigate how individual neurons (or groups of neurons) in neural networks respond to specific inputs. Among the most common neuron activation methods are activation maximisation techniques (Nguyen et al., 2016; Erhan et al., 2010). These methods search for inputs that highly activate specific neurons or layers of a neural network model to unveil trained recognition patterns; this can produce a global interpretable model and can be particularly useful for image recognition models to identify which image patches are responsible for certain neural activations.
Other:
Other approaches include emerging or ad-hoc techniques that do not fall among the traditional categories. Examples of these are Explicit Factor Model for explainable recommendations (Zhang et al., 2014), Visual Interactive Model Explorer for ML models debugging in sequential decision-making (Antar et al., 2024), approaches that exploit attention mechanisms for explaining deep neural networks (El Houda Dehimi and Tolba, 2024), and algorithmic-based (Desolda et al., 2023) or agent-based explanations (Weitz et al., 2021).
2.2. Towards Human-Centered Explainable Artificial Intelligence
Although XAI aims to address the challenges posed by AI-based systems, which, despite their high precision, often remain confined to research settings, it frequently falls short of producing usable systems that present their outcomes and explanations in a reliable, safe, and trustworthy manner. In real-world contexts, humans are generally unwilling or unable to trust black-box systems that do not provide insight into their decision-making processes (Combi et al., 2022). This lack of transparency prevents end-users from questioning AI decisions, potentially allowing harmful biases to go unnoticed (Stahl et al., 2023). In response to these challenges, a new field of study has emerged in recent years at the intersection of HCI and AI, namely HCAI (Desolda et al., 2024b). HCAI proposes a new perspective on the interaction with AI, aiming to augment rather than replace humans and their expertise (Shneiderman, 2020). HCAI systems are designed to be ethically aligned, reflect human intelligence, and consider human factors (Xu, 2019). Recent research in HCAI suggests the need for a truly human-centered process in the design of AI systems, showing how there is no “one-size-fits-all” approach when designing AI systems, as different user goals bring the need for a different level of automation and control in a system (Esposito et al., 2024). One of the milestones of HCAI is the adoption of HCI methods for the design and development of HCAI. In particular, HCAI stresses the importance of user studies in eliciting requirements and validating final systems. Human-Centered Design (HCD) is the general model adopted in HCI for the design of systems that satisfy users’ needs and expectations; it specifies that users are involved from the very beginning of the planning stage, and identifying user requirements becomes a crucial phase (Rogers et al., 2023). Various HCD-based design processes are now available, such as UCD Sprint (Larusdottir et al., 2022). HCD requires understanding who will use the system, where, and how to do what. The system is then designed by iterating a design-evaluation cycle. Being design-based on empirical knowledge of user behavior, needs, and expectations, it is possible to avoid serious mistakes and save reimplementation time to correct such mistakes. In HCAI, adopting HCD is necessary when a direct interaction between users and the system is expected (e.g., in consumer applications, educational software, or health-related support). In this context, well-designed human-centred Human-AI interfaces are essential for the success of an AI system (Shneiderman, 2022; Raees et al., 2024; Desolda et al., 2024a; Esposito et al., 2024). Furthermore, when an AI system is deployed in high-risk domains like medicine, explainability is crucial (Gerdes, 2024; the Precise4Q consortium et al., 2020; Combi et al., 2022). Therefore, a thoughtfully designed, human-centered XUI is necessary.
2.3. Explanation User Interfaces
The notion of an interface mediating between the XAI algorithm and the user was initially introduced by DARPA (Gunning and Aha, 2019). DARPA divides the XAI process into two different stages: the generation of the explanations themselves, and the presentation of such explanations through UIs. The definition of this two-stage process is shared in the academic literature. Danilevsky et al. differentiate between explanation techniques and explanation visualizations: the former involves the generation of ‘rough’ explanations (usually propounded by AI researchers), while the latter concerns the ways ‘rough’ explanations can be presented to users (Danilevsky et al., 2020). A rigorous definition of XUIs is formulated by Chromik and Butz, who define a XUI as ¡¡the sum of outputs of an XAI system that the user can directly interact with. An XUI may tap into the ML model or may use one or more explanation-generating algorithms to provide relevant insights for a particular audience¿¿ (Chromik and Butz, 2021). In general, XUIs can be designed for various forms of explanation, both local and global, and can be presented in textual or visual form, or a combination of the two  (Bove et al., 2023). XUIs can also leverage other modalities such as sound and tangible interfaces (Schneider et al., 2021a; Colley et al., 2022; Schuller et al., 2021).
According to the definition of XUIs (Chromik and Butz, 2021), two key aspects are fundamental to their development: interactivity and adaptability for different types of users. Adaptive systems change their behavior automatically, driven by context-aware mechanisms, including models of their users and of specific tasks. Adaptive systems are important because there is no ’typical’ user; there are many different users, and an individual user’s requirements can change over time (Fischer, 2023).
Interactivity plays a central role in the explanation process (Miller, 2019). Research in HCI highlights the importance of designing interfaces that allow users to explore explanations freely (Abdul et al., 2018), ask follow-up questions, and access additional details as needed (Krause et al., 2016). Achieving this level of usability in real-world scenarios requires continuous user involvement throughout the engineering process, from the initial system design to the final testing phases (Rogers et al., 2023). However, despite its emphasis in HCI, user-centred design remains relatively underexplored in XAI and XUIs, highlighting a critical gap in the field. XUIs fall into the category of interfaces designed for AI-infused systems. For these interfaces, traditional UI principles remain essential; however, the unique and dynamic behaviour of AI systems necessitates the development of new guidelines tailored to their specific characteristics (Amershi et al., 2019).
2.4. Existing Literature Reviews
Many Systematic Literature Reviews (SLRs) and surveys have explored topics closely related to our SLR, underscoring the growing relevance and urgency of advancing XAI to ensure its practical implementation in real-world scenarios. These reviews often adopt distinct perspectives, focusing on different aspects of the topic. Several studies emphasise the concept of interactivity: (Chromik and Butz, 2021) focuses on interaction design principles for XUIs, offering a structured analysis of user interactions and design guidelines. Similarly, (Bertrand et al., 2023b) categorises interactive explanation techniques based on user evaluation constructs, focusing on interactivity and its effects on user perception. (Raees et al., 2024) further advances the argument by examining interactivity in XAI systems through the lens of Human-Centred AI literature, proposing the evolution of XAI into Interactive AI. Another study (Muralidhar et al., 2023) examines the components of transparency in AI systems, aiming to enhance user interpretability and proposing strategies for designing systems that are more interpretable and explainable to users. The evaluation of XAI applications through user studies represents another important area of focus. (Rong et al., 2024) investigates how these studies are conducted, while (Al-Ansari et al., 2024) analyses the role of HCI techniques in improving XAI goals. HCI remains central to the exploration of explanation design in (Ferreira and Monteiro, 2020), whose main purpose is to investigate who the recipients of AI explanations are, the motivations for providing them, and the methods for aligning explanations with user needs. The intersection between XAI and HCI is also investigated in (Alaqsam and Sas, 2024), where the authors highlight the opportunities for HCI concerning XAI tools. Furthermore, studies like those by (Laato et al., 2022) and (Haque et al., 2022)focus on how explanations can be effectively tailored to specific user categories, particularly non-technical and end users, addressing critical gaps in accessibility and usability. Finally, (de Souza Filho et al., 2024) narrows its scope to reliance-aware explainable user interfaces, synthesising secondary studies to highlight design solutions to foster appropriate reliance in XAI-assisted decision-making.
Novelty of our SLR
Unlike other research, our review examines the literature on XUI, leveraging both algorithmic foundations and human-centered design principles. Specifically, the current research investigates three aspects: i. the contextual elements that guide the design of XUI solutions; ii. AIand XAI methodologies commonly employed in the development of XUIs; iii. the evaluation methods, driven by HCI principles, used to assess XUIs effectively.
Moreover, we also discuss the design principles that emerge from the literature. Unlike prior works that often treat these aspects in isolation, our review seeks to integrate these dimensions to identify design trends among human-centred XUIs and XAI algorithms by synthesising insights from 146 papers retrieved from IEEE, ACM, and Scopus. Our contribution lies in providing a comprehensive overview that progresses the discussion on XUIs by connecting the technical, design, and evaluation components in a unified way, addressing the need for integrated solutions in real-world XAI applications.
3. Planning and Conducting the Systematic Literature Review
We used a rigorous, reproducible methodology to perform an SLR. According to Kitchenham, conducting an SLR involves three main stages: planning, execution, and reporting (Kitchenham, 2004). This section details the first two stages, while the third one is covered in Sections 4.1, 4.2, 4.3 and 4.4.
3.1. Planning the SLR
Planning an SLR is a multi-step process that involves the following activities (Kitchenham, 2004): (1) formulating a research question; (2) defining the set of search strings; (3) electing the data sources; (4) defining the inclusion and exclusion criteria. In this section, we report the details of each activity, discussing the motivations behind every choice.
3.1.1. Formulation of the Research Questions
The main goal of our SLR is to assess the current state of research about Explanation User Interfaces and to understand how AI models, tasks, different domains, and users affect the way XUIs are designed and evaluated. With this objective in mind, we formulated the following research questions:
  1. (RQ1)
    What influences the design of XUIs?
  2. (RQ2)
    What XAI solutions (e.g., tools, frameworks, AI models) are used to develop XUIs?
  3. (RQ3)
    What solutions are used to evaluate XUIs?
  4. (RQ4)
    What can guide the design of XUIs?
3.1.2. Definition of the Search Strings
We formulated our search strings by deriving keywords from two main sources: (i) an initial assessment of some of the most highly cited papers on XUIs, and (ii) the authors’ expertise in the subject matter. The search strings were designed around three primary concepts aligned with our research questions. Each concept, along with its synonyms and variations, was included to ensure comprehensive coverage of the literature.
eXplainable Artificial Intelligence (XAI): This encompasses all techniques that provide insights into the inner workings of black-box models or involve the design of inherently interpretable (white-box) models (Guidotti et al., 2019). To capture relevant studies, we included variations of terms such as “XAI” and “eXplainable Artificial Intelligence”. The stem “expla*” was employed to include the latter and also terms like “explainability”, “explanations”, and related variations.
Human-Centered Design: High-quality user interfaces require a human-centered approach. To capture this aspect, we made sure to include keywords that could encompass terms such as “human-centred”, “human-centered”, “user-centred”, and “user-centered”, accommodating both British and American spellings. The broader terms “human” and “user” were thus used to ensure inclusivity.
Interactive User Interfaces: Since XUIs inherently involve user interfaces, relevant literature must address this aspect. Keywords like “interface” and the stem “interact*” were included to cover terms such as “interactivity”, “interaction”, and “interactive”.
The resulting final search string was defined as follows:
(”xai” OR ”expla*”) AND (”user” OR ”human”) AND (”interface” OR ”interact*”)
This formulation ensures a comprehensive yet precise retrieval of relevant studies, aligning with the scope of our SLR. In particular, our goal was to identify papers that simultaneously address all three areas — eXplainable Artificial Intelligence, Human-Centered Design, and Interactive User Interfaces — rather than focusing on each aspect in isolation. These three dimensions are further operationalized and analyzed according to the contribution types described in Section 3.2.2.
3.1.3. Selection of Data Sources
We started this SLR from scientific digital libraries. The examined libraries were ACM Digital Library [^1] https://dl.acm.org, IEEE Xplore [^2] https://ieeexplore.ieee.org, and Scopus [^3] https://www.scopus.com/. During the definition of the search string, we noticed that the same search has to be performed differently depending on the library (i.e., using a different syntax). Each library contained various options for searching content: for example, they allow searching for keywords within the article title, abstract, full text, or all of the above. We searched for the most comprehensive choice available for each digital library. All searches were conducted on the entire database due to the inherent multidisciplinarity of XUIs applications. Google Scholar was omitted, as it does not meet the inclusion criteria based on peer-reviewed publications and venue quality. Moreover, since the selected databases already cover most high-quality venues indexed by Google Scholar, its inclusion would mainly have increased the number of duplicates without improving coverage.
3.1.4. Definition of the Inclusion and Exclusion Criteria
Table 1. Inclusion and Exclusion Criteria
Criteria Inclusion Criteria Exclusion Criteria
Date Published in or after 2013 Published before 2013
Language Written in English Not written in English
Type of Publication Full and short papers Book chapters, extended abstracts, workshop proposals, posters, and demo papers
Peer Review Published in A*, A, or B conferences, or Q1/Q2 journals; C-ranked conferences or Q3 journals are carefully evaluated Published in Q4 journals, unranked conferences, or national venues
This step concerns the final selection of the relevant publications. Kitchenham distinguishes between practical inclusion and exclusion criteria, which address logistical constraints, and quality-based criteria, which are derived from the research question to refine the selection of relevant studies. Table 1 outlines the practical criteria applied in this study. Notably, we focused on publications starting from 2013 since XAI was formalized only in 2017 (Gunning and Aha, 2019), but concerns on biases and the need for explanations started earlier, as discussed in (Minsky et al., 2013). Additionally, inclusion criteria were established to ensure the selected papers aligned with the research objectives. These criteria focus on the content of the research question and require that a study:
(i) Involves a user study as part of its evaluation; (ii) Incorporates an explanation algorithm; (iii) Is based on an AI model, or, if the AI system is not explicitly defined, is designed with AI in mind (e.g., Wizard of Oz studies).
This approach ensured that only studies contributing to the advancement XAI and XUI were considered in the analysis.
3.2. Conducting the Literature Review
After planning, the literature review was conducted. As Kitchenham suggests, the review consisted of two main activities (Kitchenham, 2004): the literature review execution and the data synthesis. The subsequent subsections detail both steps.
Figure 1. PRISMA (Page et al., 2021) flow diagram depicting the identification, screening, eligibility, and inclusion process of studies in the systematic literature review.
PRISMA flow diagram depicting the identification, screening, eligibility, and inclusion process of studies in the systematic literature review.
3.2.1. Literature Review Execution
The activity was performed from June 2024 to December 2024, following the process depicted in Fig. 1, which mainly consists of two phases:
  1. (1)
    Phase 1 – Digital Library Search: We searched the identified sources using the previously described search strings.
  2. (2)
    Phase 2 – Backward and Forward Snowballing Search: We checked references and citations of the publications resulting from the previous phase and publications that cited publications from Phase 1 (Wohlin, 2014).
The initial search across the digital library yielded a total of 7084 potentially relevant publications. After a duplicate check, a dataset of 6095 publications was obtained. After filtering, using the previously defined inclusion criteria, 1023 publications were excluded as they did not meet the quality criteria, leaving a final dataset of 5072 candidates. After this first filtering step, the title and abstract were analyzed, resulting in the exclusion of a further 4658 publications, reducing the number to 414. Filtering was conducted independently by three researchers; any conflict was reported and addressed in the next step. Disagreements were resolved through discussion among the researchers, and in cases where one researcher disagreed with the other two, a majority vote was used. Inter-rater agreement on inclusion/exclusion decisions was considerable, with a Cohen’s $\kappa$ of .81.81. Through successive discussions, full inter-rater agreement was reached. The publications were then carefully analyzed by accessing their full text, narrowing down to 142 valid publications for the review. After forward and backward snowballing, an additional 148 candidates were retrieved and filtered down to 4 using the same inclusion criteria. In conclusion, the final dataset comprises a total of 146 publications.
3.2.2. Data Synthesis
We classified the types of contributions into three distinct categories:
Type 1 - UI: Publications that focus on presenting the design of an XUI tailored for a specific application.
Type 2 - Framework: Methodological papers that introduce design guidelines for XUIs or propose frameworks to support the explanation process.
Type 3 - Both: Papers that combine both aspects, presenting an XUI and proposing a set of design guidelines or a framework to support its development or evaluation.
For each paper, multiple categorical dimensions were collected for analysis, which align with established taxonomies to facilitate a structured evaluation of the results. Furthermore, for papers that reported insights for designing XUI systems—such as principles, recommendations, lessons learned, etc.—we collected design guidelines and frameworks. The knowledge extracted from the selected literature is reported in Section 4.
4. Results
The 13 dimensions identified in the literature review predominantly reflect either a human-centered or algorithmic perspective. For example, aspects related to user validation (e.g., visual mode, type of study) belong to the former category, while aspects related to the AI model (e.g., AI task, type of data) fall into the latter. The only exception is the evaluation construct dimension, which is equally applicable to both perspectives. Moreover, data from each dimension were used to answer one or more research questions. A complete description of the extracted dimensions and their corresponding research questions is presented in Table 2.
Table 2. Overview of research dimensions, associated research questions, and their descriptions.
Dimension Perspective Research Questions Description
Application domain Human-centered RQ1 Refers to the specific field or sector where the Explainable AI (XAI) system is applied, such as healthcare, finance, or education.
Type of data Algorithmic RQ1 Indicates the nature of the data used by the AI system, such as tabular, image, text, or audio.
AI Algorithmic RQ2 The specific AI model or architecture employed, such as neural networks, decision trees, or ensemble methods.
AI Task Algorithmic RQ2 Describes the goal of the AI system, such as classification, regression, clustering, or prediction.
Output Type Algorithmic RQ2 Refers to the format of the AI model’s outputs, such as probabilities, labels, and rankings.
XAI Techniques Algorithmic RQ2 The methods used to generate explanations, such as counterfactual reasoning, feature importance, and Shapley values.
Explanation Modality Human-centered RQ2 The mode of explanation delivery, such as textual, visual, natural language, or a combination of these.
Interactivity Human-centered RQ2 Indicates whether the XUI system implements interactive elements to enhance user engagement and understanding.
Visual Mode Human-centered RQ2 The type of visual representation used in the interface, such as heatmaps, trendlines, or other visual models.
Type of Study Human-centered RQ3 The methodological approach used to evaluate the Explanation User Interface, such as interviews, controlled experiments, or user observations.
Type of Users Human-centered RQ1, RQ3 The target audience for the system, such as domain experts, end-users, or AI specialists.
Number of Participants Human-centered RQ3 The sample size of users involved in studies evaluating the system.
Evaluation constructs Both RQ3 The criteria used to assess the effectiveness of the XAI system, such as trust, usability, understanding, or task performance.
4.1. RQ1. What influences the design of XUIs?
This research question investigates the factors that influence XUIs design, identifying the critical elements (e.g., constraints and considerations) that guide it. Our analysis examines the problem space from two perspectives. From a human-centred perspective, we focus on the domain of the application and the target users to whom the interface is designed. From an algorithmic perspective, we analyse the type of data utilised by the underlying AI layer.
Summary of findings
High-stakes domains (e.g., health, finance) more commonly employ XUIs, which are generally used by domain experts (e.g., physicians). Moreover, AI systems behind XUIs typically work with tabular and image data (e.g., patient data, medical scans). XUIs used in the finance domain, albeit high-stakes, are also intended for non-expert users and tend to utilize time-series and tabular data.
Table 3. User Types Dist.
Type of Users Count
Domain-experts 48
Non-experts 38
AI experts 22
Not specified 21
Table 4. App. Domains Dist.
Application Domain Count
Health 38
Finance/Economics 20
General 15
NLP 13
AI & Robotics 12
Media & Comm. 10
Education 9
Rec. Systems 7
Network 5
Mobility 4
Cybersecurity 3
Other 3
Agriculture 2
Table 5. Data Types Dist.
Type of Data Count
Tabular Data 61
Images 26
Text 19
Time Series 13
Video 6
Multimodal 4
Audio 2
Unknown 4
Generic 3
4.1.1. Application Domains
The analysis of application domains, reported in Table 5, highlights the prominence of XUIs in high-stakes fields such as health and finance (He et al., 2024; Esfahani et al., 2024; Barda et al., 2020; Wang et al., 2021; Kim et al., 2023b; Bhattacharya et al., 2024; Meza Martínez et al., 2023; Purificato et al., 2023). These domains, characterised by their critical decision-making requirements, align closely with the ones in which XAI is extremely useful (Combi et al., 2022). However, underexplored areas like mobility (Antar et al., 2024; Schneider et al., 2021a), cybersecurity (Desolda et al., 2023; Stites et al., 2021; Wu et al., 2021), and education (Shen et al., 2023; Shin et al., 2022) suggest opportunities for expanding the impact of XUIs into new contexts. The number of domains reinforces the broad applicability of XUIs solutions across various sectors.
4.1.2. Type of Users
Table 5 illustrates the distribution of user types targeted by XUI designs. The results reveal a predominant focus on domain experts, reflecting the necessity for explainability tools in decision-critical tasks within specialised industries such as healthcare and finance (Wysocki et al., 2023; Bhattacharya et al., 2023; Ferrario et al., 2020). Non-experts constitute the second-largest group, highlighting efforts to democratise AI and make it accessible to broader audiences (Wang et al., 2021; Bove et al., 2022). In contrast, interfaces targeting AI experts are less common (Shin et al., 2022; Antar et al., 2024), possibly due to their familiarity with AI systems. The substantial proportion of studies labeled as “Not specified” emphasises the need for clearer identification of user personas in future research. Furthermore, user studies often utilised platforms such as Prolific [^4] https://www.prolific.com and Amazon Mechanical Turk [^5] https://www.mturk.com, indicating a reliance on online participant recruitment for evaluation purposes (Dominguez et al., 2020; Guo et al., 2022; Alqaraawi et al., 2020; Mucha et al., 2021; Guesmi et al., 2022).
4.1.3. Data Types
The distribution of data types in XUI designs is presented in Table 5. Tabular and image data prevail, reflecting the prevalence of structured datasets and the importance of images in the most-explored domain (i.e., medicine) (Hwang et al., 2022b; Gromowski et al., 2020). This also highlights the importance of visual representation in explainability. Other types, like text (Riveiro and Thill, 2021) and time series (Upasane et al., 2021), indicate a growing interest in different data modalities. However, the relatively smaller focus on multimodal and media-based data (Wang et al., 2022) reveals potential for future exploration, particularly in complex scenarios requiring the integration of multiple data types.
4.1.4. Interplay Among Application Domain, User Type, and Data Type
Further insights from Figure 2 confirm earlier observations about the strong relation between user types and application domains. The health domain emerges as strongly associated with domain experts and both tabular data and images, reflecting the importance of precise and trustworthy explanations in high-stakes contexts such as medical diagnostics and healthcare decision-making (Combi et al., 2022). Here, XUIs are often designed to align with the expertise of professionals, ensuring that these tools support accuracy, trust, and informed decisions (Okolo et al., 2024; Zhang et al., 2024; Naiseh et al., 2023; Lamy et al., 2019).
Figure 2. Heatmap showing the relationship between user types and application domains, illustrating how different user categories are distributed across various domains
Relationship between user types and application domains. The figure illustrates how different user categories are distributed across various domains
Similarly, the finance and economics domain demonstrates a dual focus, catering to both domain experts (Purificato et al., 2023; Füßl et al., 2024), such as financial analysts, and non-experts, such as general users of financial services (Bove et al., 2022; Chromik, 2021; Bertrand et al., 2023a). Most of the time, XUI systems in this domain focus on time series and tabular data. This reflects the need for XUIs that provides clear, interpretable insights for specialists while ensuring accessibility and usability for broader audiences. The media and communication domain displays a balanced representation of general users and domain experts (without a strong focus on a specific data type) (Anderson et al., 2020; Wu et al., 2022; Khanna et al., 2022), underscoring the necessity of designing interfaces that make AI-driven explanations accessible to diverse audiences, including professionals and general users. However, AI and robotic systems are predominantly associated with AI and domain experts (Piorkowski et al., 2023; Baniecki et al., 2024; Nourani et al., 2022; Mishra et al., 2022). This trend aligns with the technical expertise required to develop, understand, and evaluate such systems. For general-purpose applications, the findings indicate a noticeable emphasis on AI experts, suggesting that such interfaces often prioritise technical depth and flexibility. These systems are frequently foundational tools or frameworks designed to be adaptable across various domains, appealing to users with advanced AI knowledge who can tailor them to specific use cases.
4.2. RQ2. What XAI solutions (e.g., tools, frameworks, AI models) are used to develop XUIs?
Figure 3. Heatmap showing the comparison between AI models and XAI techniques, illustrating the most frequent usage of XAI techniques with AI models identified in the analyzed literature.
This research question has a more algorithmic-centric perspective and examines the connection between the algorithms used to generate explanations and how these are conveyed through XUIs. To structure the analysis, we utilised the taxonomy proposed by Guidotti et al. (2019), as outlined in Section 2.1, which categorises XAI techniques based on their methodological characteristics and their application to AI models. Additionally, we consider the framework described by Gunning and Aha (2019) to classify explanation modalities (i.e., how explanations are conveyed to users through different media). In this context, the term solution is intentionally used in a broad sense, referring to any XUI component or approach—consistent with the DARPA XAI definition—that mediates the interaction between the user and the explanation provided by an XAI system (See Section 2.3).
Summary of findings
The most common XAI solutions in XUIs are feature importance, counterfactual explanations, and Shapley values, while the most common AI models are neural networks, ensembles, and transparent models. Some XAI solutions are more present for certain AI models (e.g., neural activation techniques are used with neural networks, explanation decision trees are used with reinforcement learning models). Most XUIs use visual-based explanations, most commonly employing heatmaps, barcharts, and trend lines. Textual explanations are also widely used. XUI systems often consist of interactive UI.
4.2.1. Relationships Between AI Systems and Explainability Techniques
Table 7 summarises the AI algorithms employed in the surveyed studies. Neural networks prevail over the other techniques, followed by ensemble methods and white-box models. These results highlight the prevalence of complex, opaque models like neural networks, which require XAI techniques to enhance interpretability. The distribution of XAI methodologies, as shown in Table 7, reveals that feature importance methods are the most frequently used, followed by counterfactual explanations and Shapley values. These techniques primarily aim to provide localised insights into individual model predictions. The dominance of neural networks and the frequent use of these explanation techniques underscore the need for some form of interpretability in such complex systems. Feature importance, counterfactual explanations, and Shapley values are valuable tools to bridge the gap between black-box models and actionable, user-friendly explanations. As emphasised by Miller, effective explanations should focus on relevance and comprehensibility rather than exhaustively detailing model structure (Miller, 2019). It should still be noted that, although the literature employs such techniques to provide some form of explanation for the black-box model, this has a polarising effect as some argue that this is, in fact, not the best course of action, instead preferring white-box models when possible (Rudin, 2019).
Table 6. AI Algorithm
AI algorithm Count
Neural Network 51
Ensemble 17
White Box 12
Probabilistic 7
Math 5
Rule-based 5
Reinforcement Learning 5
Agnostic 3
Other Black-box 4
Fuzzy 1
Unknown 13
Table 7. Explainable AI Techniques
Explanation Techniques Count
Features Importance 32
Counterfactual/Exemplars 28
Shapley Values 27
Exemplars/Prototype-based 19
Decision Rules 14
Saliency Mask 12
Neurons Activation 7
Decision Trees 4
Partial Dependence Plot 4
Sensitivity Analysis 1
Other 12
Some less common techniques, such as salient masks (Hwang et al., 2022b; Kadir et al., 2023), neuron activation (Wang et al., 2023), decision trees (Khanna et al., 2022), and partial dependency plots, suggest specialised applications in XUIs. These methods often target unique requirements or particular user groups, emphasizing the diverse needs within the XUIs design space.
Regarding the relationship between XAI techniques and AI algorithms, several recurring patterns can be observed. Feature importance is prominently linked to neural networks (Lee et al., 2020; Hernandez-Bocanegra and Ziegler, 2021; Röhrl et al., 2023; Lee et al., 2023; Heimerl et al., 2022; Mishra et al., 2022; Shin et al., 2022), ensembles (Malandri et al., 2023; Bhattacharya et al., 2024; He et al., 2024), and probabilistic methods (Riveiro and Thill, 2021), reflecting its adaptability across diverse AI models. Counterfactual explanations are broadly applied, particularly in ensemble (Malandri et al., 2023; Ma et al., 2022; Bove et al., 2023; Cau et al., 2023) and neural network settings (Sovrano and Vitali, 2021; Esfahani et al., 2024; Zhang et al., 2024; Hao et al., 2023). Shapley values are closely associated with neural networks (Lu et al., 2024; Kim et al., 2023c; Meza Martínez et al., 2023; Wang et al., 2022; He et al., 2024) and probabilistic models (Malandri et al., 2023; Purificato et al., 2023; Cheng et al., 2022), leveraging their utility in quantifying feature contributions in complex systems. Techniques like neuron activation are nearly exclusive to neural networks (Hwang et al., 2022b; Wang et al., 2023; Nourani et al., 2020; Ahn et al., 2024; Wu et al., 2021) due to their specific relevance to such architectures. Decision rules appear with rule-based (Piorkowski et al., 2023; Sevastjanova et al., 2021), ensemble (Bhattacharya et al., 2024; Wang et al., 2019) and Neural Network (Ming et al., 2019; Meza Martínez et al., 2023; Krause et al., 2017), while decision trees are used along with reinforcement learning techniques (Khanna et al., 2022; Kridalukmana et al., 2022).
Overall, the choice of XAI techniques often depends on the type of AI algorithm being explained. While some XAI techniques are specifically designed for particular model architectures, such as neural activation analysis for neural networks (Wang et al., 2023; Hwang et al., 2022b; Nourani et al., 2020), or analyzing decision trees in reinforcement learning (Kridalukmana et al., 2022; Khanna et al., 2022). Rule extraction techniques, although associated with rule-based models (Piorkowski et al., 2023; Sevastjanova et al., 2021), can be effectively applied to a wide variety of predictive models to generate interpretable insights, regardless of the underlying algorithm (Sabbatini, 2025; Gromowski et al., 2020; Meza Martínez et al., 2023; Metta et al., 2021; Ming et al., 2019). Model-agnostic methods, on the other hand, include feature importance, Shapley values, and counterfactual examples. These models are widely adopted due to their adaptability across various contexts and model architectures (Stites et al., 2021; Riveiro and Thill, 2021; Malandri et al., 2023; Okolo et al., 2024).
4.2.2. Explanation Modalities
The presentation of explanations to users is crucial in determining the effectiveness and usability of XUIs. Table 10 categorises the explanation modalities observed in the surveyed studies. Visual explanations dominate, likely due to their intuitive and accessible nature. Formats such as graphs, heatmaps, and diagrams translate complex AI reasoning into interpretable insights for users. Textual explanations are also widely adopted, reflecting the value of concise descriptions for clarity. Natural language explanations, audio, and tangible formats are comparatively underutilised, potentially reflecting their more specialised use cases or higher development costs. Video explanations are rare, possibly due to the time-intensive nature of their production and the immediate interactivity required in many XUIs. Figure 4 details the visualization types employed in XUIs. Heatmaps stand out as the most frequently used visualization, likely due to their ability to quickly convey patterns or feature attributions and their widespread use in image explanations in general. Bar charts and trend lines are common, reflecting their utility in comparative and temporal data analysis. Hybrid approaches that integrate textual and visual elements seem to be preferred in the literature, suggesting that combining modalities may provide the most effective explanations by complementing graphical clarity with textual context.
Figure 4. Frequency of visualisation techniques used in explanation user interfaces. Heatmaps, bar charts, and trend lines are the most frequently used. A wide range of other formats, such as scatter plots, histograms, Sankey diagrams, and tree structures, are also represented, reflecting the diversity of visual explanation strategies.
Frequency of visualisation techniques used in XUIs. The bar chart presents the types of visualisations employed to convey explanations, with heatmaps, bar charts, and trend lines being the most frequently used. A wide range of other formats,such as scatter plots, histograms, Sankey diagrams, and tree structures, are also represented, reflecting the diversity of visual explanation strategies.
4.2.3. Interactivity
Interactivity is a fundamental aspect of XUIs, influencing how users engage with and interpret AI explanations. Across the surveyed studies, we identified 95 interactive UIs and 47 non-interactive ones.
Interactive XUIs allow users to manipulate explanation parameters, explore different perspectives, or request additional details, potentially enhancing comprehension and trust (Weitz et al., 2021; Gromowski et al., 2020). These interfaces often employ dynamic visualization techniques, adjustable feature importance rankings, or interactive counterfactual explanations, enabling users to tailor explanations to their needs. Conversely, non-interactive XUIs present static explanations without user input (Riveiro and Thill, 2021; Upasane et al., 2021). While these may still provide valuable insights, they often lack adaptability, potentially limiting their effectiveness in complex decision-making scenarios. The prevalence of interactive systems in the dataset suggests a growing emphasis on user engagement in XAI research, highlighting the importance of designing explanations that support exploratory and user-driven interactions.
4.3. RQ3. What solutions are used to evaluate XUIs?
This research question is characterized from a human-centered perspective, since it explores the relationship between the user, the type of user study conducted, and the evaluation constructs (e.g., usability, trust) employed, bridging the Human-Computer Interaction perspective with the algorithmic approach to offer a comprehensive understanding of the XUIs environment. In the development of XUIs, evaluating how users interact with and perceive these systems is critical to ensuring their effectiveness, usability, and trustworthiness (Abdul et al., 2018). The way users are evaluated and the constructs observed play a central role in understanding whether the explanations provided by an AI system align with user needs and expectations. By examining study methodologies and key performance metrics, researchers can identify strengths and gaps in XUI design, ensuring that these systems are technically robust and user-centered. This analysis is essential for advancing the field of XAI and fostering greater adoption of AI systems in real-world applications where transparency and trust are crucial.
Summary of findings
The most common evaluation methods used with XUIs are controlled experiments (typically used with domain experts and non-experts), followed by interviews (often used with domain experts, capturing qualitative data) and usablity studies. Domain experts are typically involved in evaluating XUIs in high-stakes domain (e.g., health and finance), while AI experts are also involved in usability studies and interactive feedback and co-design sessions. Evaluation studies often measure trust, usability, workload, perceived effectiveness, and satisfaction. Constructs such as helpfulness are typically measured with domain experts, while AI experts are generally involved in studies measuring task performance, usability, and confidence.
Table 8. Expl. Modalities
Modality Count
Visual 76
Text 50
Natural lang. 6
Audio 3
Tangible 2
Video 1
Other 2
Table 9. Type of User study
User Study Count
Controlled Exp. 53
Interview 28
User Observation 18
Usability Study 15
Survey 10
Focus-Group 4
Interactive session 4
Wizard-of-Oz 2
UI inspection 2
Table 10. Top 14 Eval. Constructs
Eval. Constructs Count
Trust 33
Understandability 30
Usability 24
Satisfaction 20
Usefulness 15
Workload 15
Perceived Effec. 14
Task Perform. 14
Transparency 8
Perc. Quality 8
Helpfulness 8
Confidence 7
Ease of use 6
Explainability 6
4.3.1. Evaluation Methods and Constructs in XUI Research Table 10
categorises the various study methods used to evaluate XUIs. Controlled experiments are the most frequent ones, highlighting the importance of rigorous, systematic evaluation under controlled conditions. This method allows researchers to isolate specific factors and measure their impact on user interaction with XUIs. Additionally, interviews and usability studies are frequently utilised, highlighting the focus on user perspectives and ensuring that the interfaces remain practical and accessible.
Moving onto the evaluation constructs, illustrated in Table 10, trust stands out as central, indicating that fostering user confidence in AI systems is a key priority for XUI developers. Usability and workload are close behind, emphasizing the importance of ensuring the interface is easy to use and does not overwhelm the user. Perceived effectiveness and satisfaction suggest an interest in understanding users’ perceptions of how well the system supports their needs. Table 11 shows the most popular evaluation constructs. Collectively, these constructs underscore the importance of designing XUIs that not only enhance explainability but also encourage practical adoption in real-world applications. XUIs should, in fact, be designed to inspire trust, remain user-friendly, and avoid overwhelming users with excessive complexity.
It is worth noting that, despite the importance of transparency in explainability research, it has only been explicitly assessed in eight studies, suggesting that current research in XUIs should place greater emphasis on incorporating transparency constructs into evaluations.
Table 11. Papers that report user studies evaluating Trust, Usability, and Workload
Evaluation Constructs N. Papers References
Trust 28 (Weitz et al., 2021; Kim et al., 2023a; Wysocki et al., 2023; Guo et al., 2022; Khurana et al., 2021; Lee et al., 2020; Panigutti et al., 2023; Lu et al., 2024; Karran et al., 2022; Bhattacharya et al., 2023; Meza Martínez et al., 2023; Hernandez-Bocanegra and Ziegler, 2021; Bhattacharya et al., 2024; Kim et al., 2023d; Röhrl et al., 2023; Guesmi et al., 2022; Joshi et al., 2024; Lombardi et al., 2024; Naiseh et al., 2023; Bahel et al., 2024; Guesmi et al., 2024; Lee et al., 2023; Perlmutter et al., 2024; Bertrand et al., 2023a; Park et al., 2022; Zytek et al., 2022; Purificato et al., 2023; Heimerl et al., 2022; Nimmo et al., 2024; Lei et al., 2024; Desolda et al., 2023; Stites et al., 2021)
Usability 24 (Fujiwara et al., 2020; Wu et al., 2022; Füßl et al., 2024; Khurana et al., 2021; Wang et al., 2021; Huang et al., 2022; Cheng et al., 2021; Kim et al., 2023b; Szymanski et al., 2024; Hohman et al., 2019; Guesmi et al., 2024; Wang et al., 2022; Ming et al., 2019; Sevastjanova et al., 2021; Park et al., 2022; Zhang et al., 2024; El-Zanfaly et al., 2023; Sciascio et al., 2017; Purificato et al., 2023; Hao et al., 2023; Yuan et al., 2024b; Antar et al., 2024; Szymanski et al., 2021)
Workload 15 (Fujiwara et al., 2020; Piorkowski et al., 2023; Dominguez et al., 2020; Sanneman and Shah, 2022a; Lee et al., 2020; Wang et al., 2023; Karran et al., 2022; Khanna et al., 2022; Anderson et al., 2020; El-Zanfaly et al., 2023; Kridalukmana et al., 2022; Sciascio et al., 2017; Nakao et al., 2022; Heimerl et al., 2022; Yuan et al., 2024a)
Figure 5. Methods used to evaluate explanation interfaces across different user types. The heatmap shows the distribution of evaluation methods, categorised by the type of user involved (e.g., AI experts, domain experts, non-experts).
Methods used to evaluate explanation interfaces across different user types. The heatmap shows the distribution of evaluation methods such as controlled experiments, interviews, surveys, and usability studies, categorised by the type of user involved (e.g., AI experts, domain experts, non-experts). Controlled experiments and interviews are the most common methods, particularly among non-experts and domain experts, reflecting a strong emphasis on both empirical validation and qualitative feedback in XUIs evaluation.
4.3.2. Tailoring XUI Evaluation Methods to User Groups
Figure 5 represents the distribution of user types across various study methodologies used in the evaluation of XUIs. The figure provides some insights into how different user groups are targeted and about the methodological approaches used to assess their interaction with and understanding of the interface. The choice of study method is influenced by the target user group: controlled experiments are favoured for domain experts and lay users due to their rigorous structure. Interviews, however, are particularly prominent in studies involving domain experts, as they can usually capture qualitative insights. Domain experts are typically involved in interviews for the application of XUI technologies in the high stakes domains such as Health (Hwang et al., 2022b; Kim et al., 2023a, c; Bhattacharya et al., 2023, 2024; Naiseh et al., 2023; Okolo et al., 2024; Larasati et al., 2023; Lindvall et al., 2021; Zhang et al., 2024; Cheng et al., 2022) and Finance and economics, (Füßl et al., 2024; Purificato et al., 2023) but also in other domain such as Education (Lu et al., 2024), Natural Language Processing (Wang et al., 2022) and weather forecasting (Kim et al., 2023d). Methods such as interactive feedback sessions, co-design, usability studies, and, notably, surveys are also employed to evaluate AI experts in various domains, such as health (Hohman et al., 2019) associated with Artificial Intelligence (Baniecki et al., 2024), education (Shin et al., 2022; Yuan et al., 2024b) and finance and economics (Hohman et al., 2019; Meza Martínez et al., 2023). While these approaches are often associated with assessing non-expert users, their application to AI experts suggests an interest in gathering qualitative and usability-focused insights from technically proficient individuals.
Notably, certain study types, such as Wizard-of-Oz (Hernandez-Bocanegra and Ziegler, 2023; Joshi et al., 2024) experiments, exhibit a high proportion of unspecified user roles, indicating that these evaluations often prioritise AI behaviour simulation over explicitly defined user categories.
4.3.3. User group and Evaluation Constructs
Figure 6. Reported evaluation criteria by user type. Figure shows which user groups were associated with various evaluation criteria used in assessing explanation interfaces. Criteria such as trust, understandability, usability, and perceived effectiveness were evaluated across all user types, highlighting the multidimensional nature of user-centered evaluation in explainable systems. Notably, transparency was not evaluated in studies involving AI experts.
Reported evaluation criteria by user type. The stacked bar chart illustrates which user groups (AI experts, domain experts, non-experts, and unspecified) were associated with various evaluation criteria used in assessing explanation interfaces. Criteria such as trust, understandability, usability, and perceived effectiveness were evaluated across all user types, highlighting the multidimensional nature of user-centered evaluation in explainable systems. Notably, transparency was not evaluated in studies involving AI experts.
The relationship between the user type involved in the study and the constructs assessed provides insight into the key aspects that XUI designers prioritise when evaluating an application’s suitability for real-world adoption. Figure 6 reveals several key trends. Domain experts are the predominant group assessing helpfulness (Hwang et al., 2022b; Nourani et al., 2022; Liu et al., 2022; Zytek et al., 2022); this might indicate that XUIs are designed with expert decision-making support in mind, where explainability is expected to enhance task performance rather than merely provide technical insights. Since domain experts are the end-users of AI systems in high-stakes applications (e.g., healthcare, finance), their evaluation of helpfulness is crucial in determining whether explanations effectively aid decision-making rather than just offering transparency. Interestingly, AI experts are almost absent in studies assessing transparency, despite their technical expertise. This suggests that transparency is being evaluated primarily from a user-facing perspective rather than from an algorithmic interpretability standpoint. This construct frequently appears in the Conversational AI domain (Khurana et al., 2021; Hernandez-Bocanegra and Ziegler, 2023) and recommendation systems (Hernandez-Bocanegra and Ziegler, 2021; Guesmi et al., 2022; Lee et al., 2023; Park et al., 2022), highlighting its importance in ensuring domain experts and lay-users understand system decisions and interactions in these applications. AI experts are more involved in studies that assess task performance, usability, and confidence: task performance is typically assessed in applications within computer science-specific domains, such as networking (Fujiwara et al., 2020; Xuan et al., 2022) and machine learning (Piorkowski et al., 2023). Usability studies involving AI experts span a broad range of domains, including mobility (Antar et al., 2024), education (Yuan et al., 2024b), economics (Purificato et al., 2023; Hohman et al., 2019), and agnostic applications (Cheng et al., 2021; Hao et al., 2023), reflecting a general need to ensure the usefulness of XUI applications. The confidence constructs (Karran et al., 2022; Baniecki et al., 2024; Zhou et al., 2021), measured significantly less frequently compared to the others, indicates a concern for how well the system supports trust in its outputs, ensuring that explanations are not only technically accurate but also foster reliability and acceptance in real-world applications.
4.4. RQ4. What can guide the design of XUIs?
Several papers in our review propose guidelines for designing XUIs. These emerged either as lessons learned from evaluating XUIs within a user study or as indications of a broader framework. From the manual analysis of the retrieved papers, 80 of them present a XUI and indicate clear design guidelines, while 10 solely present a theoretical or conceptual framework that can guide the design of XUIs.
Table 12 reports the list of papers that present design guidelines. We grouped the papers that propose an XUI according to the AI model used by the system. This categorization is meant to provide readers with a quick reference for the design of XUIs for systems that use a specific technology (i.e., Neural networks, Ensemble methods, Reinforcement Learning, Transparent Models, Generic AI models, Other models). Also, the list of papers that only present a framework for designing XUIs is reported.
Summary of findings
XUIs should be designed using a user-centered approach, accommodating users’ mental models and favouring transparency and trust. Providing interactivity and control of explanations, providing multiple levels of visualizations, together with contextual information to better understand them, can help in this regard, as well as allowing users to personalize explanations based on their profile. Finally, popular frameworks can guide the design of XUIs by defining how explanations should be shaped (e.g., which questions they should answer), what values a XUI system should possess (e.g., transparency, engagement, control on interactivity), how users may be involved in the design process, and how to align user goals with system explanations.
Table 12. Papers that report design guidelines divided by AI model
AI model Num. papers References
Neural networks 29 (Weitz et al., 2021; Hwang et al., 2022b; Fujiwara et al., 2020; Krause et al., 2017; Dominguez et al., 2020; Kim et al., 2023a; Wang et al., 2021; Lee et al., 2020; Huang et al., 2022; Malandri et al., 2023; Wang et al., 2023; Kim et al., 2023c; Nourani et al., 2022; Meza Martínez et al., 2023; Hernandez-Bocanegra and Ziegler, 2021; Kim et al., 2023d; Guesmi et al., 2024; Lee et al., 2023; Wang et al., 2022; Esfahani et al., 2024; Lindvall et al., 2021; Zhang et al., 2024; Hao et al., 2023; Heimerl et al., 2022; Chotisarn et al., 2023; Gehrmann et al., 2019; Mishra et al., 2022; Ming et al., 2019; Karran et al., 2022)
Ensemble methods 9 (Wysocki et al., 2023; Chromik et al., 2021; Wang et al., 2019; Cheng et al., 2022; Bhattacharya et al., 2024; Bove et al., 2023; Ma et al., 2022; Kalamaras et al., 2019; Cau et al., 2023)
Reinforcement Learning 3 (Khanna et al., 2022; Park et al., 2022; Mishra et al., 2022)
Transparent models 5 (Guo et al., 2022; Nakao et al., 2022; Bhattacharya et al., 2023; Hohman et al., 2019; Lee et al., 2020)
Generic/Model-agnostic models 30 (Ferrario et al., 2020; Sanneman and Shah, 2022a; Cabour et al., 2023; Su et al., 2020; Naiseh et al., 2024; Mosca and Such, 2021; Fiok et al., 2022; Raees et al., 2024; Liu et al., 2022; Laato et al., 2022; Meyer and Zhu, 2024; Sanneman and Shah, 2022b; Ferreira and Monteiro, 2020; Simkute et al., 2022; Wu et al., 2022; Füßl et al., 2024; Khurana et al., 2021; Jansen et al., 2024; Cheng et al., 2021; Szymanski et al., 2024; Schneider et al., 2021b; Collaris and Van Wijk, 2020; Kim et al., 2023e; Naiseh et al., 2023; Okolo et al., 2024; Bahel et al., 2024; Larasati et al., 2023; Zytek et al., 2022; Szymanski et al., 2021; Lei et al., 2024)
Other AI models 4 (Riveiro and Thill, 2021; Bove et al., 2022; Piorkowski et al., 2023; Desolda et al., 2023)
Frameworks only 10 (Cabour et al., 2023; Su et al., 2020; Naiseh et al., 2024; Fiok et al., 2022; Meyer and Zhu, 2024; Sanneman and Shah, 2022b; Simkute et al., 2022; Raees et al., 2024; Liao et al., 2020; Pieters, 2011)
4.4.1. Core design principles for XUIs
Our review of XUI design guidelines across various AI models—including Neural networks, Ensemble methods, Reinforcement learning, Transparent models, Generic approaches, and other AI models—reveals several cross-cutting principles. These core principles serve as a foundation for effective XUI design, even as model-specific requirements require tailored adaptations. In the following, we discuss these principles, also illustrating them within different AI approaches.
User-Centered Design
At the heart of any effective XUI is a design that is grounded in the user’s needs, context, and expertise. At a high level, explanations in XUIs should use clear, jargon-free language that adapts to the user’s context, reducing cognitive load and making AI behaviour more accessible (Khurana et al., 2021; Jansen et al., 2024; Okolo et al., 2024; Szymanski et al., 2024). User-centered frameworks such as (Barda et al., 2020) should be leveraged to define the purpose, content, and presentation of explanations in contexts such as clinical decision support (Hwang et al., 2022b). Ensemble-based XUIs emphasise the need to account for users’ mental models and to minimise cognitive load (Wysocki et al., 2023; Chromik et al., 2021), while generic systems advocate for tailoring explanations to both novice and expert users (Khurana et al., 2021; Jansen et al., 2024). Particular effort should be put into this aspect by researchers (e.g., using participatory design, conducting usability studies), as often solutions that appear effective to developers are not ideal to end-users (Guo et al., 2022).
Interactivity and Control
Effective XUIs empower users by supporting active engagement with the system. Supporting interactive exploration was found useful for users, for example, allowing selection, filtering, juxtaposition, and smooth transitions between views (Wang et al., 2021; Hao et al., 2023). Moreover, features such as what—if exploration, guided navigation, and real-time updates were found to enhance user engagement and comprehension (Collaris and Van Wijk, 2020; Liu et al., 2022). Interaction is also a core requirement for white-box models to iteratively refine the understanding of explanations (Guo et al., 2022; Nakao et al., 2022). Regarding reinforcement learning systems, interaction modalities such as questioning the model and navigating through explanation spaces help users clarify why actions were taken, increasing trust and understanding of the system (Mishra et al., 2022).
Transparency and Trust
User trust is fostered by a high level of transparency of the AI’s decision-making process. To build trust, the potential and limitations of the AI system should be properly conveyed by ensuring that users receive accurate cues about model confidence and reliability (Okolo et al., 2024; Zytek et al., 2022). With black-box models such as neural networks, transparency can be achieved by displaying detailed insights, such as model uncertainties, confidence scores, and error metrics (Lee et al., 2023; Malandri et al., 2023; Nourani et al., 2022; Zhang et al., 2024). Offering descriptive statistics and summaries of model performance can allow users to gauge overall system reliability (Lee et al., 2023; Malandri et al., 2023). White-box systems, in particular, stress that explanations should faithfully represent how the underlying model functions, including both local and global insights (Nakao et al., 2022; Guo et al., 2022), possibly to increase understanding of the risk behind a model’s decision (Bhattacharya et al., 2023). To increase trust, explanations should use an adequate language that is either accessible and avoids technical jargon (Okolo et al., 2024) or is tailored to the user’s domain and cultural context (Szymanski et al., 2024). Explanations should also resemble human reasoning and the way of explaining (Kim et al., 2023e). Finally, special care should also be taken in the selection of features to be addressed in explanations, as not every feature may be understandable and meaningful to the end user (Desolda et al., 2023).
Multi-level Visualizations
Effective visual representations in XUIs are essential for helping users understand complex AI behavior. Offering visual summaries through multi-level adjacent explanation visualizations was found to improve user confidence (Karran et al., 2022). Transitions and animations between different visualizations can help navigate the XUI, improve understanding of the underlying AI model, and increase learning engagement and enjoyment (Wang et al., 2021). In reinforcement learning, an effective visualization of state spaces should provide both high-level overviews and detailed insights into specific states (Mishra et al., 2022). XUIs can also benefit from visualizations such as color-coded risk factors, feature importance charts, and graphical comparisons that clarify both instance-specific and overall model behavior, also by showing counterfactuals and similar instances (Bhattacharya et al., 2023; Hohman et al., 2019; Lee et al., 2020).
Contextual Information
Placing explanations together with additional contextual information enhances their relevance and interpretability to domain-expert users, e.g., physicians (Hwang et al., 2022b; Cheng et al., 2022), AI engineers (Fujiwara et al., 2020; Bhattacharya et al., 2024), etc. Moreover, the depth of information in explanations should be adaptable to the user’s expertise, ensuring that both detailed and summary information are available without overwhelming the user (Wang et al., 2019; Naiseh et al., 2023; Lei et al., 2024).
Personalization and Adaptability of Explanations
The results from the literature stress the need for culturally sensitive and user-tailored content, ensuring that explanations are accessible and meaningful for diverse audiences (Okolo et al., 2024; Bahel et al., 2024). Designs should tailor content to the user’s background, cognitive abilities, and cultural context. This includes strategies such as incremental disclosure of information and the use of both textual and visual elements to match different learning styles (Kim et al., 2023e; Naiseh et al., 2023).
4.4.2. Frameworks
The ten frameworks resulting from the literature review cover heterogeneous aspects of the design of XUIs and may be grouped into three categories: (1) Interaction and Content Design Frameworks: These frameworks focus on the artifact (i.e., the explanation itself) and provide operational guidance on mapping user inquiries to specific interface components and explanation techniques. (2) Axiological and Relational Design Frameworks: These frameworks focus on human values and the relationship between the user and the AI, guiding the system’s high-level ethical positioning and prioritizing trust/confidence calibration, user engagement, and situational awareness. (3) Cognitive and Learning Alignment Frameworks: These frameworks focus on cognitive aspects of the user, drawing upon cognitive psychology and learning theories to align the XUI with human information processing capabilities, literacy levels, and learning goals.
Interaction and Content Design Frameworks
In (Meyer and Zhu, 2024), Meyer and Zhu examined 28 XUIs and linked specific UI components to explanation types. While “Why” and “What-if” questions are commonly addressed, tools for “How” or “Input” explanations are scarce, with elements like sliders, buttons, and icons playing key explanatory roles. Liao et al. (2020) identified gaps between XAI research and design practice by interviewing 20 practitioners, resulting in an XAI question bank based on prototypical user inquiries. These questions are shaped by user context and mapped to suitable explanation techniques, such as global or counterfactual methods. Su et al. (2020) analyzed 40 mobile health apps and found that users want AI features—like recommendations or predictions—explained in clear, engaging ways. The study emphasises aligning explanations with user literacy and providing meaningful feedback loops to refine AI behavior.
Axiological and Relational Design Frameworks
The Value Sensitive Design methodology (Naiseh et al., 2024) emphasises embedding human values—such as transparency, autonomy, and inclusivity—into explainable AI systems. Similarly, Raees et al. (2024) advocate for user engagement and transparency to enhance trust in interactive AI. Sanneman and Shah (2022b) proposed SAFE-AI, a framework for enhancing user situation awareness by addressing what AI did, why it did so, and what it will do. It also includes trust calibration and workload management to optimise user-AI interaction. Pieters (2011) distinguishes between explanation-for-confidence and explanation-for-trust, showing how explanation goals shape user understanding; the author argues that explanation depth must match its intent: too little detail may hinder trust, while too much may lower confidence.
Cognitive and Pedagogical Alignment Frameworks
In education, Fiok et al. (2022) advocate for participatory, user-centered design in XAI tools. Their work highlights the importance of grounding design in psychological theory and leveraging HCI expertise to align technical systems with learners’ needs and trust-building goals. Simkute et al. (2022) presented design principles for educational XUIs based on cognitive psychology; effective learning support includes techniques such as cognitive forcing, analogies, contrastive reasoning, and interactive feedback to promote skill development. The Abstracted Explanation Space (Cabour et al., 2023) offers a structured framework that links user goals with appropriate explanation types, covering perception, comprehension, and projection. This framework advocates cross-disciplinary collaboration and helps designers tailor explanations to avoid cognitive overload and enhance relevance.
5. Discussion
Figure 7. Distribution of explanation modalities across different application domains, data types, and types of users. Visual explanations dominate across most contexts, while other modalities, such as natural language, non-verbal, or tangible explanations, vary depending on the domain and target user group. The visualization reveals distinct preferences and trends in how explanations are tailored based on the nature of the data and the expertise of the intended users.
Distribution of explanation modalities across different application domains, data types, and types of users. The chart highlights how visual explanations dominate across most contexts, while other modalities—such as natural language, non-verbal, or tangible explanations—vary depending on the domain and target user group. This visualization reveals distinct preferences and trends in how explanations are tailored based on the nature of the data and the expertise of the intended users.
This section discusses the results by relating the dimensions analyzed in our RQs. dditionally, to make our discussion actionable, we present a platform to support practitioners in building effective XUIs: HERMES. We examine how each dimension interacts across all RQs to identify recurring patterns. This approach allows us to distill insights that can support practitioners in navigating the broader landscape of the findings.
Visual and textual explanations are the primary means of conveying explanations, as shown in Figure 7, where the explanation modality is shown relative to its use within different domains, types of data, and types of users. This preference is particularly evident when working with tabular data and images, where graphical representations can make patterns and relationships more apparent. However, visual explanations are not limited to these data types; they are also applied to text and time series, while textual explanations remain common for tabular data. It is also interesting to note how in the mobility domain (Schneider et al., 2021a), audio is present.
Certain explanation methods are widely used across domains. In healthcare, SHAP (Barda et al., 2020; Wysocki et al., 2023; Kim et al., 2023b; Cheng et al., 2022; Zhou et al., 2021), feature importance (Lee et al., 2020; Röhrl et al., 2023; Lei et al., 2024; Mishra et al., 2022), and counterfactual explanations (Lamy et al., 2019; Zhang et al., 2024) are frequently employed to interpret model decisions. These techniques are also common in finance and economics, where counterfactual reasoning (Ferrario et al., 2020; Ma et al., 2022; Sovrano and Vitali, 2021; Hohman et al., 2019; Bove et al., 2021; Esfahani et al., 2024; Cau et al., 2023) and SHAP values (Meza Martínez et al., 2023; Chromik, 2021; Purificato et al., 2023) help practitioners assess risk and model fairness. Since these methods are well-suited for tabular data, they naturally find application in domains where structured datasets are prevalent.
The choice of explanation technique often depends on the audience. SHAP (Malandri et al., 2023; Ma et al., 2022; Chromik, 2021; He et al., 2024; Jansen et al., 2024) explanations tend to be preferred when working with non-expert users due to their intuitive visualization of feature contributions. However, exemplars i.e., concrete instances that illustrate a model’s behavior, and salient masks, which highlight the most influential input regions, are also commonly used to aid understanding (Huang et al., 2022; Wang and An, 2021; Yang et al., 2020; Larasati et al., 2023; Nakao et al., 2022). When designing user studies for XUIs, the choice of method should align with the target user group. Interviews are most commonly used for domain experts, particularly in high-stakes fields like healthcare and finance, where qualitative insights are crucial for understanding decision-making needs. In contrast, surveys and usability studies are more prevalent among AI experts, emphasizing practicality and efficiency in technical contexts. Regarding evaluation constructs, domain experts in high-stakes applications tend to prioritise helpfulness (Hwang et al., 2022a; Nourani et al., 2022; Liu et al., 2022; Zytek et al., 2022) over transparency (Sevastjanova et al., 2021). Their primary concern is whether an explanation supports their decision-making process rather than simply revealing model internals. For XUIs designers, this means that explanations should be crafted to disclose system behaviour and integrate seamlessly into expert workflows, ensuring they provide meaningful, actionable insights. The SLR presented in this paper has drawn on works from multiple disciplines to gather knowledge from diverse expertise fields. These vary substantially according to several factors, such as the application domain, the technologies underlying an AI system, and the type of users. Therefore, accessing these findings and integrating them into a specific design context would be cumbersome and require extensive knowledge of different fields (e.g., a designer should be knowledgeable about AI technologies). To make the findings more accessible, we developed a web platform called HERMES, which is described in the following.
5.1. Navigating the Results: Introducing the HERMES Platform
[…]
Figure 8. The HERMES Framework with the 6 dimensions observed in our analysis.
The HERMES Framework with the six dimensions observed in our analysis
Figure 9. A screenshot of the HERMES platform showing guidelines for specific defined criteria: (AI Model = “Neural network”, AI Task = “Binary classification”, Application Domain = “Health”, and Type of Users = “Domain-experts”.
A screenshot of the HERMES platform showing guidelines for specific defined criteria: (AI Model = “Neural network”, AI Task = “Binary classification”, Application Domain = “Health”, and Type of Users = “Domain-experts”.
HERMES (a framework for Human-cEnteRed developMent of Explainable user interfaceS) is a web platform designed to support practitioners in building effective XUIs. It offers a structured collection of design guidelines derived from the results of the SLR, presented as interactive cards that synthesize insights from the referenced papers. Each card includes a concise description of a guideline, relevant tags that reflect key dimensions from the associated paper, and a direct link to the original source. Users can explore the guidelines in two complementary ways: by browsing the full collection of cards, or by filtering them through up to six entry points (Fig. 8), which correspond to six of the thirteen dimensions of this survey (see Table 2): (1) Application Domain, (2) User type, (3) AI model, (4) AI task, (5) XAI model, and (6) Explanation modality. HERMES was developed using Jekyll [^6] https://jekyllrb.com/ and enables designers to either align their XUIs with an existing use context (by analyzing the above dimensions) or to explore potential design directions under specific project constraints, fostering both conformity and creativity in the design process. Not all the input dimensions must be defined by the designer. Moreover, the output of the platform also includes data about the techniques to employ for the evaluation of the XUI (i.e., the Type of Study to conduct and the evaluation constructs to assess). The platform is available at the following address: http://espositoandrea.github.io/hermes-mirror/. A screenshot of the web platform is reported in Fig. 9.
5.2. Use case: a medical application
Let us suppose that a medical technology company is developing a XUI for an existing AI-based system designed to assist clinicians in diagnosing cervical cancer using MRI scans. The existing software that currently helps clinicians is a neural network-based system that determines whether a patient’s blood test suggests that the clinical criteria for diabetes are met. However, the AI model only gives the clinician a binary outcome (yes/no), without providing an explanation. The development team is thus tasked to implement an XAI model and visualise the outputs for clinicians, who need transparent and interpretable AI outputs to support their diagnostic decisions.
To guide the design of this XUI, the development team uses HERMES to determine the most suitable explanation and UI design strategies by setting four out of the six entry points as project constraints: Health as the application domain, Neural Network as the AI model, Image classification (cancerous vs. non-cancerous MRI scans) as the AI task, and Domain Expert as the user type, as the interface is targeted to domain experts (i.e., clinicians). Since there are no other design constraints, the other two entry points (XAI model and Explanation modality) are left empty – thus becoming suggestions from HERMES. Based on the provided parameters, HERMES suggests explanation methods that have been effective in similar medical applications. The system highlights several XAI techniques as suitable, including Counter-exemplars/factual and Decision Rules, and Natural Language, Text, and Visual explanations as explanation modalities. HERMES also identifies the design guideline Provide Contextual Information, which states that placing explanations alongside additional contextual information enhances their relevance and interpretability for domain-expert users, such as physicians, while tailoring the depth of explanations to the user’s needs. To support this recommendation, HERMES provides references to relevant academic sources, including a specific paper by Cheng et al. (2022). In the paper, the developers can find more implementation details and access a visual example to guide the design of the new XUI. To ensure the XUI meets clinicians’ needs, HERMES suggests that user studies for designing and evaluating such interfaces should include interviews and user observation studies. This may be interpreted as a recommendation to conduct interviews with clinicians early in the design process, and then use observational studies to iteratively test and refine the XUI, following a human-centered design approach (ISO, 2019b). Using HERMES, the design team efficiently selects an adequate explanation approach tailored to expert clinicians in a medical domain. The platform informs the selection of XAI models and explanation modalities, and also suggests a human-centered approach to involve end-users both in the design and evaluation of the XUI. This structured methodology can indeed accelerate the design process, reducing trial-and-error efforts and ensuring that the final interface aligns with best practices found in the academic literature.
6. Future Challenges
Among the future challenges identified by this SLR, a key issue is the limited adoption of co-design practices and focus group methodologies in the development of HCXAI systems. These approaches are essential to strengthen a real co-creation dynamic between users and XAI systems, ensuring that user needs, mental models, and expectations are meaningfully integrated into the design process. Similarly, the collection of quantitative data and the execution of usability studies involving both AI and domain experts are critical for increasing the practical applicability and reliability of these tools. Recent research acknowledges the growing role of Large Language Models in XAI. Lei et al. suggest that with the rising adoption of LLMs, the demand for XAI among non-expert users will increase. They propose a set of design principles for XUIs (Lei et al., 2024). LLMs are also beginning to be incorporated into Explainable Conversational Interfaces. For example, Joshi et al. conducted a Wizard of Oz user study comparing two versions of a vacation planning chatbot—one with low explainability and one with high explainability—involving 60 participants. Their study finds that providing explanations enhances trust and acceptance of the LLM-based system (Joshi et al., 2024), LLMs represent both a novel challenge and a promising opportunity for the development of XUIs. They are reshaping user-explanation interactions and advancing the potential for more conversational, human-like explanations, as envisioned by Miller (Miller, 2019). However, LLMs themselves remain opaque, prone to hallucinations, and can undermine user trust due to their lack of transparency and controllability. Addressing these matters is key for the future of effective and responsible XUIs design.
A significant future challenge for the HERMES platform lies in the development and validation of more specific and context-aware design guidelines, particularly those that intersect multiple design dimensions. Currently, many guidelines remain broad or generic, limiting their applicability to complex or highly specialised use cases. Advancing HERMES will therefore require the formulation of more granular recommendations tailored to particular combinations of application domains, user types, AI tasks, and explanation modalities. To achieve this, integrating crowd-sourcing mechanisms could play a crucial role. By involving a diverse community of practitioners and researchers, it would be possible to both enrich the guideline repository with real-world insights and validate existing guidelines through collective evaluation. This participatory approach could significantly enhance the relevance, reliability, and adaptability of the platform, making it a more robust decision-support tool for the design of XUIs across varied contexts.
7. Limitations and Threats To Validity
In general, several threats to validity can undermine the results of a Systematic Literature Review. In the following section, we report the most common ones, detailing how we mitigated them. Selection Bias: This occurs when the studies included in the review are not representative of the entire population of studies on the topic. This has been mitigated by: i) manually reviewing the publications to ensure their compliance with the SLR goal, and ii) performing two phases, i.e., search on digital libraries and snowballing. However, our search focused only on full publications; therefore, we acknowledge that this SLR does not cover demos, posters, and working papers presented at workshops. Publication Bias: This occurs when studies that show statistically significant results are more likely to be published than studies that do not. This aspect has been mitigated by manually reading those publications that do not report any results but only a technical solution with preliminary results. Besides the generic inclusion criteria, their relevance for our SLR is considered, for example, the number of citations and the novelty of the solution. Time Lag Bias: This occurs when the review does not include all relevant studies because they were published after the review was conducted. In this case, we can safely assume that this threat is not so evident in our study since the SLR was performed 5 months before its submission. Publication Quality: This occurs when studies of poor quality are included in the review. To mitigate this aspect, we defined inclusion criteria on the quality of the venue of the publication, leaving to a manual evaluation of the authors of this SLR the inclusion of publications that appeared in venues of lower quality. Grey Literature was not Considered: This includes preprints and tools, such as commercial tools and platforms, that were not published in academic venues, despite their popularity or relevance to our RQs.
The organization and navigation criteria adopted in this SLR should not be interpreted as exhaustive or definitive since they come from the authors’ design choices. Alternative classification schemes or additional dimensions may be equally valid and could offer complementary insights, particularly for domain-specific or practice-oriented use cases. Moreover, the effectiveness and appropriateness of these navigation criteria have not yet been empirically evaluated with end users. To mitigate this limitation and enable stakeholders to customize their navigation and obtain insights into our results, we provide access to HERMES. Additionally, while we acknowledge the importance of user-based validation of HERMES, we consider it as a supporting and exploratory artifact within the broader scope of this work—which synthesizes and structures existing knowledge on XUIs—rather than as a standalone, fully validated system. HERMES was in fact designed to complement this contribution by offering an interactive and actionable way to navigate the results of the SLR, facilitating exploration by researchers, designers, and practitioners. A comprehensive user-centered evaluation of HERMES (in particular assessing usability, usefulness, and impact on research and design practices) is therefore left as an important direction for future work. Finally, findings were reported by clusters to maintain contextual specificity and avoid overgeneralization across multiple domain contexts. This approach indeed limits the interpretational boundaries of the results since cross-domain transferability would require dedicated user studies, thus strengthening the methodological rigor of the SLR and preventing misleading conclusions.
8. Conclusions
This study explores the design and evaluation of XUI. While XUIs are widely recognised as a fundamental component of the explanation process, their implementation still presents significant research challenges. We approached this problem from a multidimensional perspective. On one hand, we examined HCI requirements, analyzing user types and evaluation constructs derived from user studies. On the other hand, we investigated algorithmic aspects, bridging these two dimensions through the interface layer. Our analysis led to the development of guidelines for designers, combining both top-down and bottom-up approaches. The top-down perspective involved reviewing studies that propose directives and frameworks, while the bottom-up approach focused on extracting design principles and requirements from case studies. These guidelines aim to close the gap between the usability of explainable interfaces and approaches that focus solely on algorithmic transparency. By providing a comprehensive view of the ecosystem in which XUIs are designed and implemented, this study helps identify key challenges in creating explainable processes that comply with emerging regulations. Ultimately, our findings contribute to making XAI not only a research concept but a practical and deployable solution across real-world applications. In fact, in this work, we also introduced HERMES, a platform for guiding practitioners in the design and evaluation of XUIs. Although user studies to evaluate and validate the platform are needed, with this platform, we attempt to bridge the gap that exists between results coming from the academic literature and their actual application by designers and developers.
Authors Contribution Statement